Tickets to the county fair for four adults and five children cost $33. An adults ticket costs $1.50 more than a childs ticket. Find the cost of an adults ticket and write an equation.
Answer by mananth(16949)

child ticket $x
Adult (x+1.50)
5x+4(x+1.50)=33
5x+4x+6=33
9x=33-6
9x=27
/9
x=3 child ticket
Adult ticket = (3+1.50)= $4.50